OpenStudy (anonymous):

prove that there exists no prime such that p-q=97,if p and q are prime

OpenStudy (anonymous):

when p and q are both not equal to 2 then, p and q are both odd numbers and difference of two odd number is an even number so one of them has to be even. Now, when p=2 then 2-q=97 ==> q=-95 [not prime] when q=2 then p-2=97 ===>p=99[not prime] Thus, there exists no prime such that p-q=97,if p and q are prime

OpenStudy (anonymous):

great thanks I just have to formalize it

OpenStudy (anonymous):

What can I say now.. I can just say that @sauravshakya is right...

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

Ha ha ha ha..